OpenStudy (anonymous):

Solve for x. 2x + 16 = 3(x - 9)

OpenStudy (jack1):

2x + 16 = 3(x - 9) so 2x + 16 = 3x - 27 now move the x's to one side and the whole numbers to the other 16 + 27 = 3x - 2x can you solve from here?

OpenStudy (anonymous):

We have : \[2x+16=3x-27\] Then : \[2x-3x=-27-16\] Then : \[-x=-43\] Then : \[x=43\]